OpenStudy (anonymous):

In a noninjury, noncontact skid on icy pavement on an empty road, a car spins 1.75 revolutions while it skids to a halt. It was initially moving at 15.0m/s, and because of the ice it was able to decelerate at a rate of only 1.50m/s^2. Viewed from above, the car spun clockwise. Determine its average angular velocity as it spun and slid to a halt. Answer is 1.10 rad/s, down. I need to know how to work the problem please.

OpenStudy (anonymous):

Knowing how many revolutions it spined, you can get how many radians, knowing its initial velocity, and the aceleration you can get how long it took for the car to stop. Then, to calculate the average velocity (angular or usual) you divide the space that the object passed (radians or meters), by the time it took to pass.

OpenStudy (anonymous):

0 =15 - 1.5*t t=10s (theta f - theta i )/10s = omega average 1.75 rev = 3.5pi rad clockwise from above is negative direction (-3.5pi -0)/10 = omega avg = -1.099 rad./sec